2/20/23- Arrived on site and spoke with facility manager. Unit located in the basement in the back, was at temp upon arrival. Checked over unit and did not see anything that was concerning. Put unit into defrost manually multiple times to ensure the compressor kicked back on. Condenser fan motor on start up was slow and under amping. Replaced condenser fan motor and is rejecting heat nicely. Ran and came down in temp and satisfied.
Arrived at location on 8/1/22 - Removed racks and blower wheel guard. The screws and plate were caked on with hard grease. Gas stove chip away to removed the bolts. Removed old motor. Installed new. Reassembled all. Turned on. Blower works in cook and cool down modes. Oven heats ok. Back in service.
Arrived at location on 7/15/21 - checked unit found paddle switch not opening. For final rinse. Pulled cover and switch is broken. Put back together and let Joe know of problems. Told him if there not washing dishes to turn dishmachine off to preserve rinse water. Unit is operational but requires part. Returned on 7/20/21 - removed cover for switch. Removed box and took switch out. Wired in new switch and reinstalled all. Adjusted arm nut. unit is operational job complete.
Arrived on site to check Vulcan range, before turning gas on performed a quick once-over of unit. Preceded and turn gas on and turned on Pilots. Checked operation of top burners and found right front pilot corroded and blocked. checked other Pilots, okay. Top burner operation okay, left rear pilot needs to be replaced. Checked griddle operation, okay. Lit Oven Pilots and checked operation. Both ovens and checked operation okay. Could not locate any gas leaks. Will order new pilot tip for right front burner. Adjusted both oven doors. Range in service. 11/16/18 Turned off gas to unit. removed front panel to get at pilot tube. removed burners. removed right front pilot tube. removed pilot tip and replaced with new. reassembled and lit pilots. right front pilot barely lit and then went out. tried to adjust pilot flame but nothing happens. removed pilot tube and have no clogs in pilot tube or tip. found that pilot adjuster valve orfice is clogged. cleaned and replaced. reassembled and lit p
Arrived at location to check steam table, left weld needs a new element because wire and prong to element is corroded and broke off, thus not letting weld heat to temp. Infinite switch on middle weld is going bad which is causing middle weld not to heat up to proper temp so middle needs new infinite switch and the last weld to the right needs a infinite switch because it's broken and no knob can't be put on unit, also needs knobs as well. Need to order all parts QWP (expedited freight) as per Larry the director of the rehab center. Called eagle metal master and spoke with Donna for part numbers and prices. Will order parts through shop and return for install. 9/13/18 Arrived at location, replaced all three of the infinite switches, replaced old and broken element from left well and put on all three of the knobs on infinite switches.. tested unit, let it get up to temp, unit is heating. Unit back in service
Found from left #1 & 2 are leaking at o-ring & sockets behind the mixing bowls on CEC Coffee Machine. Also #5 leaks to #2 English toffee valve leaks by. Tried cleaning valve seat but still drips by. Need to order o-rings and one valve. Will return. QWP (expedited freight) 10/20/17 Removed bowl sockets, o-rings & valve. Replaced with new. Reassembled. Powered on. Unit works with no leaks. In service.
Checked elements all pull 8.6 amps. Put water in kettle and ran. kettle did have good vacuum on it when cold. Unit wont heat past # 7 . Tried adjusting . Temp. control is flaky . Recommended replacing temp. control and thermistor. Gave prices, customer okay to order. 4/21/17 Installed temp control and thermistor. Unit heats a bit then shuts off. Hi limit trips out. Changed hi limit. Unit has a rolling boil @ #7. Unit in use.
Replaced blower motor and blower wheel per quote.
Turned Frymaster fryer on to 350. Burners came on. Heated to 350. Cycling at 350. Let drop in temp. cycled back on at 344. Let cycle a few times. Made adjustment to t-stat. Brought cycle on temp up. Knob has no numbers on it. Marked with a screwdriver until new knob arrives. Temp degrees go to 400F. Chef said have knob shipped to this location.
Adjusted doors on Convection Oven. Now closing okay. Chef verified. Back in service.